Do It! Review 20-2

Snow Cap Springs produces and sells water filtration systems for homeowners. Information regarding its three models is shown below.

Basic

Basic Plus

Premium

Total

Units sold 930 345 225 1,500
Selling price $254 $396 $802
Variable cost $199 $292 $413

The company’s total fixed costs to produce the filtration systems are $128,007.
Determine the sales mix as a function of units sold for the three products. (Round answers to 0 decimal places, e.g. 12.)

Basic

Basic Plus

Premium

The sales mix as a function of units sold

%

%

%
Determine the weighted-average unit contribution margin. (Round answer to 2 decimal places, e.g. 10.25.)
Weighted-average unit contribution margin $

Determine the total number of units that the company must produce to break even.
Break-even point in units

units
Determine the number of units of each model that the company must produce to break even.
Basic

units
Basic Plus

units
Premium

units
Answer #1
Basic Basic plus Premium
Sales mix (930/1500) = 62% (345/1500) = 23% (225/1500) = 15%

.

Weighted average unit contribution margin

= [(254-199)*62%]+[(396-292)*23%]+[(802-413)*15%]

= 116.37

.

Breakeven units = Fixed cost/WACM

= 128,007/116.37

= 1100 units.

.

Basic (1100*62%) 682
Basic plus (1100*23%) 253
Premium (1100*15%) 165